top of page

DEPOT REPORT SATURDAY 22nd JUNE

Updated: Jun 25

VLO, 2 new potential volunteers attended for their Induction. Gary Blackburn & Callum Lilley. Gary stayed on and joined the team and Callum had to leave but has committed to return.


55009. The 2nd nameplate was reunited with the loco on 'B' side, unfortunately the supplier only sent 2 sets of numbers so we are awaiting those to complete the full return to 55009 and now the '13' has gone hopefully the gremlins will leave as well.


55015. A very busy day with lots of cable tracing going on. This started in No 1 nose end with 6 unidentified but possibly from the boiler controls, they were securely taped up. In No 2 end cubicle No 4 a relay was found to be sticking, it was removed, fixed, tested and replaced. A systematic check on all relays is being done, again, as they were done a good while back.

Enough progress was made to be able to close the BIS and the CCB. On closing the BIS the No 2 end fire alarm bell rang continuously and the No 1 end did not ring at all. This is being investigated, possibly the ER Fire Sensors are open circuit. No 2 engine Fuel Priming Pump worked successfully but unfortunately the No 1 engine pump did not, suspect fault in a coil, investigation next week.


55019. The first of the 2 main Res tanks, 'B' side was collected from Tecforce and was installed, with the other being side having a temporary one loaned from 9 the batteries were connected, the air air system was charged from the static compressor following the fitting of the 2 main res tanks today and no leaks found. A fabricator was on site on Thursday and fitted the 2 new Nose End Footsteps, these were dressed and painted with primer ahead of the painting team completing the job.

The batteries and the AVR's were reconnected, the BIS was closed and all was ok and now ready for testing next week.


Many thanks to all who attended, a great days work.


Pics by Nigel Julian and Paul Beard.






450 views0 comments

Recent Posts

See All

Comments


bottom of page